[ please help me solve this question. 2x-4y=-1 and y=x-1. they are to the same problem. I can't find the answer. please help me ] NOTE: apparently the problem asks you to find the intersection of the two lines, if any.
Found 2 solutions by ewatrrr, TimothyLamb:
Answer by ewatrrr(24785)

 
Hi,
2x-4y=-1 &
y=x-1
finding the point of Intersection or the 'solution' to this system of EQs
2x - 4y = -1
2x - 4(x-1) = -1 |Substituting for y y+=+x-1+
-2x +4 = -1
-2x = -5
x = (5/2) and y = (3/2) (x-1)
Pt (5/2, 3/2) is the solution for this system of EQs (the lines intersect at that point)


Answer by TimothyLamb(4379)

---
2x - 4y = -1
y = x - 1
---
put the system of linear equations into standard form
---
2x - 4y = -1
x - y = 1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
the point of intersection of the two lines is:
x= 2.5
y= 1.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php